The length of a rectangle is 4 in. more than twice its width. If the perimeter of the rectangle is 56 in., find the width of the rectangle.
Answer: The width is ______

Answer by checkley75(3666)

L=2W+4
P=2L+2W THUS
56=2(2W+4)+2W
56=4W+8+2W
56=6W+8
6W=56-8
6W=48
W=48/6
W=8 IS THE WIDTH
PROOF
56=2(2*8+4)+2*8
56=2(16+4)+16
56=2*20=16
56=40+16
56=56